Susan is a single and has the following situation for the year: Susan's income of $80,000; dividend income of $20,000; interest income of $2,000; short-term capital gain of $8,000 and long-term capital gain of 14,000. She also paid $1,000 on interest charges on her credit card. Her other total exemptions and itemized deductions is 22,000; these amounts will be deducted from her gross income to determine her taxable income. If she is files as a single individual, what is Susan's marginal tax rate? Use the individual tax rate provided below.

Individual Tax Rates: Single Individual

Taxable Income

You Pay This Amount on the Base of the Bracket

Plus This Percentage on the Excess over the Base (Marginal Rate)

Average Tax Rate at the Top of Bracket

Up to $8925

$0

10.0%

10.0%

$8925-36350

$892.50

15.0

13.8

$36250-87850

$4991.25

25.0

20.4

$87850-183250

$17891.25

28.0

24.3

183250-398350

$44603.25

33.0

29.0

$398350-400000

$115586.25

35.0

29.0

Over $400000

$116163.75

39.6

39.6

  1. 10.0%

  2. 20.0%

  3. 25.0%

  4. 28.0%

  5. None of the above
Answer #1

Answer:

Correct answer is:

28.0%

Explanation:

Susan's income = $80,000

Dividend income = $20,000

Interest income = $2,000

Short-term capital gain = $8,000

Long-term capital gain = $14,000

Gross Taxable income = 80000 + 20000 + 2000 + 8000 + 14000 = $124,000

Less: Itemized deduction = $22,000

Taxable income = 124000 - 22000 = $102,000

Long term capital gain = $14,000

Taxable as ordinary income = 102000 - 14000 = $88,000

Tax bracket is $87850-183250 and the tax you pay is $17891.25 plus 28 Percentage on the Excess over the Base (Marginal Rate)

Susan's marginal tax rate is 28%

Hence option D is correct and other options A, B, C and E are incorrect.